a fishing enthusiast buys 50 maggots and 20 worms for $1.10 and her mother buys 30 maggots and 40 worms for $1.50.Find the cost of one maggot and one worm
Answer by nerdybill(7384)

a fishing enthusiast buys 50 maggots and 20 worms for $1.10 and her mother buys 30 maggots and 40 worms for $1.50.Find the cost of one maggot and one worm
.
Let x = cost of on maggot
and y = cost of one worm
.
50x + 20y = 1.10
30x + 40y = 1.50
.
-100x - 40y = -2.20
30x + 40y = 1.50
---------------------
-70x = -0.7
x = -0.7/(-70)
x = 0.01 (one maggot cost 1 cent)
.
cost of worm:
50x + 20y = 1.10
50(0.01) + 20y = 1.10
.5 + 20y = 1.10
20y = .6
y = .6/20
y = 0.03 (one worm cost 3 cents)